Solve the system:
(x+3)^2+(y+2)^2=17
x+y=-2
Enter your solutions as an ordered pair. Write "none" if there are no solutions.

Answer by ewatrrr(24785)


Hi
x+y=-2 0r y =-x-2
(x+3)^2+(y+2)^2=17
(x+3)^2+(-x)^2=17
2x^2 + 6x - 8 = 0
x^2 + 3x - 4 = 0
(x+4)(x-1) = 0, x = -4,1
(-4, 2) and (1, -3) are the solutions for this System